A is an aqueous solution of H2SO4 of concentration 0.05 moldm3,B is put A intoa burette pipette 25cm3 of B into a conical flask and add two drops of methyl orange indicator



What is the concentration in moldm3 of NaOH in B given the equation for reaction H2SO4+2NaOH



What is the percentage purity of the NaOH



What is the percentage of the impurity present

1
Expert's answer
2022-04-05T15:53:03-0400

Since the volume of H2SO4 is not given, let's assume it was 20.0cm3.

Moles of H2SO4 = 0.05×20÷1000= 0.001

Moles of NaOH= 0.001×2= 0.002

Concentration of NaOH= (0.002÷25)×1000= 0.08M


Percentage purity of NaOH

Mass of NaOH in 1dm3= 0.08×40= 3.2g

Assuming density of solution is 1g/cm3,

Percentage purity=( 3.2÷1000)×100

=0.32%

Percentage impurity= 100-0.32

= 99.68%